find the mass of a conical funnel z=sqrt(x^2+y^2) 0<=z<=4 if the density per unit area is p=8-z
Answer #1

Griven. Z = of 2<4 s-density 8-2 1 0 <834 ) Z=r Pwinmeterize. &= 8-8 x = X y=y ds -Zx-Zy , 1) z=f(x,y) dis X 102 tyr ✓x²+ ye